Outline of Eng Tests

  1. LBC IT 7513 - test for jumps in LBC - sidereal guiding
    • Observe field with PA=0 and PARANGLE ~ 90 deg with long ~180-sec exposures (dithering fine)
    • Use pcsclient to send a 0.2-0.3" guiding offset and see how LBC reacts
  2. LBC NS guiding test - IT 7382 - no dithering - long exposures, bright object and faint object (hedge against weather issues)
  3. Binocular MODS/LBC-R test.
    • IT 7500. LBC-R script needs to have a line added that specifies that the type of offset (async = 0, sync[default]=1). The line is of the same form as the XDIT and YDIT lines: copy and paste on of these and then edit it to be like: <SYNCOFFSET> 0, 0, 0, 0 </SYNCOFFSET>
  4. Pseudo-monocular guiding test - MODS/LBC-R, LUCI/LBC-R
    • Set of scripts for Caldwell objects @ 6-12 hrs being made
  5. Pointing MODS (Poor conditions)
  6. Pointing LUCIs (Poor conditions)
  7. ESM tests - HCG 31 (1 hour)
  8. Poor Weather Science
    • H-band Library (Poor conditions - bad seeing and thick clouds) - BinoLUCi
    • CO-bandhead LIRG merger (end of night) - BinoLUCI
-- BarryRothberg - 19 Dec 2018

OSA: D.G. Huerta

ISp: E. Solheid

Mtn Manager: R. Blizzard

ISA: B. Rothberg (Tucson)

Observers: B. Rothberg, O. Kuhn (Tucson)

-not open at handover due to clouds

UT 08:40 - reconfigured for MODS-1 (SX), LBC-R (DX) - will attempt closed dome tests if we can't open.

When I brought up mods2Disp, I noticed that the the last MODS2R mask image displayed had some horizontal trails from the slits in the left quadrant to the left edge. A charge transfer effect to investigate. Not seen on sieveSnap.

-errors on LBC -

Ran up LBC blue + red successfully (or so I thought... but the log does not show that the TURNON command was executed --- this was probably the problem), but when I tried to run an OB (2 Bias Bino Check), it hung. Log was showing many messages like the following, where the IIF_initialize failed. (We had LBCR authorized, not both, so it is odd that it reports that the IIF_initialize failed for both.)

2018/12/22 08:40:48.283331 W -                   TCS IIF_initialize failed for prime both focstation, LBC_DD_Proxy proxy
2018/12/22 08:40:48.283354 W -                   Failed setting LBC float DD items [src/lbc/lbc.c:2098]
2018/12/22 08:40:48.283585 W -                   TCS IIF_initialize failed for prime both focstation, LBC_DD_Proxy proxy
2018/12/22 08:40:48.283606 W -                   Failed setting LBC string DD items [src/lbc/lbc.c:2142]
2018/12/22 08:40:48.283821 W -                   TCS IIF_initialize failed for prime both focstation, LBC_DD_Proxy proxy
2018/12/22 08:40:48.283841 W -                   Failed setting DD item LBCOBProgress [src/lbc/lbc.c:2072]

GUI is hung and I cannot kill/start the SW:

[lbccontrol@cmu ~]$ lbckill
[lbccontrol@cmu ~]$ lbcstart
lbccontrol: Starting lbc 
LBC control software is already running!

Rebooted CMU at 09UT. Powered back on LBCR only, this time. Filter and rotator problems... Try again, same thing.

2018/12/22 09:06:30.289544 E R ROTATOR           GOYA bad communication with serial device [src/rotator/rotator.c:487]
2018/12/22 09:06:22.066861 E R FILTERS  WHEEL#1  bad communication with serial device [src/filters/filters.c:516]

Powered off HK, Powered back on, this time, LBCB only. Took 2 LBCB biases and this time the OB did not hang. 09:21 Turning on Red only now. It came up and we took 2 LBCR biases successfully. We still have nothing authorized on the right. Connected for right side and all OK.

09:35. PCS patch 20181219 has been activated. Ready to go.

09:38 David bounced PCS.

09:47. Test sidereal guiding on field at ra, dec = 05:30, 45:00. It has parangle = 92. It is close to moon & cloud.

There was a problem visualizing this field with the OT. Downloaded an Image (archive.eso.org) but it gave me the DSS image around 14:10, -17:18 and there is a Copyright at the top right edge, as if it is the corner of an image/plate. I can successfully get an image for a field at 05:17:24, 34:22:48.

09:52 dohybrid, /redonly (plenty of stars...). Converged, estimated seeing 0.9".

09:59 Play OB. This will take an 150/120-sec exposures at r/z, and do a 5pt dither.

10:03 Restart OB after insuring that PCS was bounced after patch activated.

10:09 Guiding seems to be OK, but r-band is saturated (so close to moon). Moving to another field (12h45d).

10:18 dohybrid, /redonly (ooops... should have just run dofpia)

10:23 Play OB. Found guide star, using 8.0-sec. parangle = -90 deg.

lbcr.20181222.102400.fits 150-sec sky background 43000 (stopped to decrease exptime and refocus - guide star becoming elongated). Stars on science image look round-ish, though, and no jumps seen.

10:30 dofpia. 105-sec r-band exposure started (103249). Issued SetCentroid DX 0.5 0

10:37: new r-band exposure. SetCentroid DX 1.0 0 at 10:38 (during 103700).

10:41 revert patch and repeat. Reverted patch @ 10:42 (to 20180924).

10:45. Playing OB again. At 10:46:41, issued SetCentroid DX 1.0 0 (during 104539).

Issued SetCentroid DX 2.0 0 at about 10:50:30. (104956 has some elongation)

10:54. Put 20181219 patch back & bounce PCS.

10:55 Play OB again. 10:56:55 issue SetCentroid DX 2.0 0 (during 105610, which has 3 images for each star).

10:59:45 issue SetCentroid DX 0.0 2.0 (during 105807). 11:01:33 0 1.0 during (110131).

There is only 1 image which showed 3 star images, and this one was with the new patch, so the results of this test are not so clear.

Leaving the patch in place as we go to a NS target. VR5.

11:09 Slew to VR5 copointing/collimation field. dofpia.

11:18 NSIGUI -> Set to hijack. Play OB for VR5. Think I see the object - remains in the same place but stars move from image to image.

11:33 Slew to VR copointing/collimation field. dofpia.

11:38 NSIGUI -> Set to hijack. Play OB for VR. (By mistake, I played the OB which dithered). See object - it looks round (but faint) and the star trails look pretty straight.

11:53 Slew to ET2 copointing/collimation field. dofpia.

11:59 NSIGUI -> Set to hijack. Play ET2 nodither OB. Hard to find this one.

12:14 Authorize MODS1/SX and LBCR/DX. Collimate SX near first target, NGC2775 (09:10:20, 07:02:16).

12:37 Collimated on MODS1/SX. Authorized LBCR only to slew to LBC collimation field.

12:40 Authorize MODS1/LBCR.

acqMODS --mods1 --bino MODS1_NGC2775_pa345.acq.txt

Play LBCR_NGC2775_pa345.xml (with syncoffset line added to make offsets async).

12:43 Sync preset completed. LBCR 124353 exposure started. Rotational trailing.

from log,

2018/12/22 12:42:51.087618 W R ROTATOR           no tracking trajectory supplied, rotator tracking is disabled [src/rotator/rotator.c:1890]

But 2nd exposure 124702 is OK and the log reports that a trajectory was loaded.

12:52 execMODS --bino --mods1 MODS1_NGC2775.obs.txt

NOTE: 125304 - LBCR image has lots of white (NaN) pixels. I noticed this while looking at the 20181210 data. Will IT it.

13:08. The MODS OB dithered up/down by 2". It is most likely that hte dither happened during an LBC exposure, but the exposure looks OK. And LBC is dithering during MODS exposures, and these don't seem to be seen. The two sides are offsetting independently. The only problem seems to be the first LBC exposure which is rotationally trailed.
Data are:
mods1r.20181222 (4-7) and mods1b (2-5)
and
lbcr.20181222.124353.fits[2] NGC 2775 0 0
lbcr.20181222.124702.fits[2] NGC 2775 0 0
lbcr.20181222.124955.fits[2] NGC 2775 -2 -11
lbcr.20181222.125304.fits[2] NGC 2775 -2 -11
lbcr.20181222.125608.fits[2] NGC 2775 2 11
lbcr.20181222.125921.fits[2] NGC 2775 2 11
lbcr.20181222.130219.fits[2] NGC 2775 -11 2
lbcr.20181222.130528.fits[2] NGC 2775 -11 2
lbcr.20181222.130832.fits[2] NGC 2775 11 -2
lbcr.20181222.131145.fits[2] NGC 2775 11 -2

13:18:17 - LBCR rotator trajectory failed. I replayed the OB which had a preset... cannot do that.

13:25 - Take 2 short, 30-sec, r and z band exposures using fast-OB GUI and "Current" pointing. And take a u/i image pair with MODS1, manually through the gui.

13:21 Feige67 with MODS1 (Authorize MODS1/NONE).

13:34 acqMODS --mods1 feige67.acq. GCS stuck at the step of initializing WFSCam. Restarting AZCam Server.

13:51 Starting LBC skyflats at i-SLOAN (sky brightness already 13.5). Too bright - got 1 or 2 good i-SLOAN ones, no V-BESSEL flats.

14:01 Put MODS1 and MODS2 to sleep. Turned off the LBCR.

We tested the patch for LBC guiding and it worked although it was hard to pinpoint. Our tests to 'nudge' the telescope did not show much difference before/after patch, but someone looking at the logs may be able to pull out more. Ran LBC in NS and binocular mode and with the patch installed, and images were mostly good, with the exception of rotational trailing on the first LBC exposure of the OB.

LBC electronic issues to investigate?
Topic revision: r10 - 04 Jan 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback